A wheel has a string of length 4 m wrapped round its shaft. The string is pulled with a constant force of 150 N. It is observed that when the string leaves the axle, the wheel is rotating at 3 revolutions in a second, find the moment of inertia of the wheel. 3.38 kg-m*)


1
Expert's answer
2021-02-17T11:08:08-0500
"Tr=I\alpha\\(150)(4)=(3\cdot 2\pi)I\\I=31.8\ kg\cdot m^2"